OpenStudy (anonymous):

QUICK Will Medal!!!! 4/9 divide 3/7=

OpenStudy (anonymous):

Find LCD

OpenStudy (anonymous):

I don't have a calculator ATM 😭

OpenStudy (anonymous):

Lowest Com man Denominator

OpenStudy (anonymous):

give me a sec let me figure it out

OpenStudy (anonymous):

kk who ever does it first gets the medal or just a hapy face :D

OpenStudy (anonymous):

\[\frac{\frac{4}{9}}{\frac{3}{7}} = \frac{4}{9}*\frac{7}{3} = \frac{4*7}{9*3}\]

OpenStudy (anonymous):

good job zelda

OpenStudy (anonymous):

that would equal 28 over 27

OpenStudy (mathmale):

LCD has nothing at all to do with this problem. It's a problem involving division. Please start over. "4/9 divide 3/7=" Do you mean "4/9 divided into 3/7," or "3/7 divided into 4/9?"

OpenStudy (anonymous):

whaa?

OpenStudy (anonymous):

look

OpenStudy (mathmale):

whaa is not a question. Please spell it correctly. "kk who ever does it first gets the medal or just a hapy face" It's you, Nico, who needs to solve this problem first, not your helpers.

OpenStudy (anonymous):

i already answered the question so ima close this now

OpenStudy (mathmale):

the problem you have posted is one of division, not of addition or subtraction.